Suppose you need $4000 in 3 years for a trip to Europe. How much do you have to deposit into an account each month if the annual interest rate is 3.5% compounded monthly in order to meet your goal? Round your answer up to the next dollar. Express your answer with no dollar sign and no decimal point
Answer by Boreal(15235)

4000=m[(1+ i/t)^Nt-1 ]divided by i/t
4000=m(1+.035/12)^36/.035/12
multiply by denominator (.035/12)
11.66667=m(1+0.0029166667)^36-1)=m(.11054) without rounding, divide by 0.11054
106
This can be checked roughly by multiplying by 36 months, which gives $3816, not far from $4000, and with 3.5% interest would be reasonable to expect this amount to reach $4000 in three years.